Q: solve the pair of linear equations by the method of equating the coefficients :- x/a=y/b.....(1) and ax+by=a^2+b^2.....(2) .

Grade:10

1 Answers

Rinkoo Gupta
9 years ago
x/a=y/b
=>x=ay/b
substitute this value in second eq
a(ay/b)+by=a^2+b^2
a^2y+b^2y=b(a^2+b^2)
(a^2+b^2)y=b(a^2+b^2)
y=b
x=a(b)/b=a
{x=a,y=b} Ans.
Thanks & Regards
Rinkoo Gupta
askIITians Faculty